Required information [The following information applies to the questions displayed below.] Morganton Company makes one product and it provided the following information to help prepare the master budget a. The budgeted selling price per unit is $70. Budgeted unit sales for June, July, August, and September are 9,400, 25,000, 27,000, and 28,000 units, respectively. All sales are on credit. b. Thirty percent of credit sales are collected in the month of the sale and 70% in the following month. c. The ending finished goods inventory equals 20% of the following month's unit sales. d. The ending raw materials inventory equals 10% of the following month's raw materials production needs. Each unit of finished goods requires 4 pounds of raw materials. The raw materials cost $2.50 per pound. e. Twenty percent of raw materials purchases are paid for in the month of purchase and 80% in the following month. f. The direct labor wage rate is $15 per hour. Each unit of finished goods requires two direct labor-hours. g. The variable selling and administrative expense per unit sold is $2.00. The fixed selling and administrative expense per month is $64,000
9. If 108,800 pounds of raw materials are needed to meet production in August, what is the estimated raw materials inventory balance at the end of July? Raw material inventory balance
10. What is the total estimated direct labor cost for July assuming the direct labor workforce is adjusted to match the hours required to produce the forecasted number of units produced? Total direct labor cost
11. If we assume that there is no fixed manufacturing overhead and the variable manufacturing overhead is $10 per direct labor-hour what is the estimated unit product cost? (Round your answer to 2 decimal places.) Unit product cost
12. If we assume that there is no fixed manufacturing overhead and the variable manufacturing overhead is $10 per direct labor-hour what is the estimated finished goods inventory balance at the end of July? Ending finished goods inventory
Answer #1

Unit produced = 25000+(27000*20%)-(25000*20%) = 25400 Units

9) raw material inventory balance = 108800*10%*2.5 = 27200

10) Direct labor cost = 25400*2*15 = $762000

11) Unit product cost = (15*2+2.5*4+10*2) = 60 per unit

12) Ending finished goods inventory = 27000*20%*60 = $324000